# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### 05 Apr 2024 v2

@@ 5.2. Active Learning Algorithms
Figure 5. Top: The Active Learning Framework. Bottom: The relationship and correspondence between each component of Active Learning and those of CAT.
The idea and the pattern of active learning are similar to that of CAT, inspiring a general framework. To be specific, in CAT systems, the CDM serves as the learning model, the question selection algorithm serves as the data selection strategy, and the examinee’s question-taking behavior can be regarded as data annotation. The goal is to select as few samples (questions) as possible to effectively train the learning model (estimate the proficiency parameters in CDM). This framework allows for a model-agnostic perspective in designing CAT solutions, as it doesn’t require the consideration of the CDM’s specific implementation details. For example, Bi et al. (2020) propose a model-agnostic adaptive testing framework along these lines. Its core idea is that the change (i.e., the gradient in machine learning) in the proficiency estimate parameter, upon observing an examinee’s response to a question,
(8) $j_{t+1}=\arg\max_{q_{j}\in\mathcal{Q}}\mathbb{E}_{y_{j}}\left\|\nabla_{\theta}L(X_{1:t}\cup\{(q_{j},y_{j})\}|\theta)\right\|.$
It calculates the expected norm of each resulting gradient vector, taking into account the unknown response label $y_{j}$ during the selection process. The intuition behind this framework is that it prefers questions that are likely to most influence the proficiency estimation (i.e., have the greatest impact on its parameters), regardless of the resulting label. The active learning technique successfully alleviates the model-specific issues in CAT.
Discussion 3.
The above heuristic rules-based approaches (i.e., statistical algorithms and Active Learning algorithms) cannot be improved by learning from large-scale examinee response data (Ghosh and Lan, 2021; Li, 2020). To overcome this limitations and enable scalable personalized testing, data-driven approaches have gained prominence. These approaches using machine learning and deep learning can achieve better performance. They leverage techniques such as Reinforcement Learning (Section 5.3) and Meta Learning (Section 5.4) to automatically make suitable question selection by learning patterns from large-scale examinee response data.
5.3. Reinforcement Learning Methods
@@
Reinforcement learning (RL), a subfield of machine learning, is a powerful approach that enables an agent to learn how to make optimal decisions automatically (Sutton and Barto, 2018). It has been successfully applied in various domains, including robotics, autonomous vehicles, education, and healthcare (Arulkumaran et al., 2017; Kober et al., 2013). In RL, an agent interacts with an environment and receives feedback in the form of rewards or penalties based on its actions. The goal is to learn a policy $\pi$, which can maximize the long-term cumulative reward. As shown in Figure 6, the policy can be learned by exploring the environment and learning from its consequences of actions. In essence, researchers in CAT utilize RL methodologies to address a question: Can the selection algorithm (policy) be automatically learned and optimized from data or student interactions, thus circumventing the necessity for expert intervention? Markov Decision Process Formulation. The interaction between agent and environment can be viewed as a Markov Decision Process (MDP) (Feinberg and Shwartz, 2012). Specifically, at each step, the agent observes current environment’s state ($s$), and interacts with the environment by selecting its actions ($a$). Simultaneously, the agent receives a reward ($r$) from these interactions, influencing or changing the current state of the environment. The objective is to select a best sequence of actions, resulting in the highest cumulative reward ($\sum r$). Therefore, most RL problems are formally described as estimating the optimality of the agent’s behavior in a given state (value-based methods (Sutton and Barto, 2018)) or the optimality of the action policy itself (policy-based methods (Agarwal et al., 2021)) or the hybrid approaches (Mnih et al., 2016). We next formulate the CAT problem as an MDP. Here are the definitions of the key elements in RL within an adaptive learning/testing system and the overall RL framework of CAT is in Figure 6:
Figure 6. The overall Reinforcement Learning framework of CAT. The objective is to optimize the selection algorithm $\pi$ (i.e., policy) by exploring the large-scale examinee response data (i.e., environment).
@@
  • State: A state $s_{t}\in S$ represents the current condition or situation at each test step $t$. It captures relevant information about examinee and the CAT system. Generally, the state includes the examinee’s previous response sequence (or a latent vector to represent the current proficiency estimate (Li et al., 2020)) and the candidate questions in the question bank (Zhuang et al., 2022b; Nurakhmetov, 2019): $s_{t}=(\{q_{1},y_{1},...,q_{t},y_{t}\},\mathcal{Q})$.
@@
  • Reward: A reward $r$ is a scalar feedback that the CAT receives after selecting a question for the examinee at each step. To achieve CAT’s goal (i.e., $\min{\|\hat{\theta}^{T}-\theta_{0}\|}$ in Definition 1), the reward function can be defined as the accuracy of proficiency estimation at each step [^1] As the true value is often unobtainable, it is commonly derived through simulation experiments. (Ma et al., 2023; Nurakhmetov, 2019; Wang et al., 2023a), i.e., $\|\hat{\theta}^{t}-\theta_{0}\|$, or the performance prediction loss of $\hat{\theta}^{t}$ on the held-out response data (Zhuang et al., 2022b; Shin and Bulut, 2022). This reward signal is pivotal in guiding the policy $\pi$ to select the best-fitting question that can reduce the estimation error.
Recently, with the advancements in deep learning, an increasing number of studies are leveraging Deep Reinforcement Learning (DRL) to tackle the MDP problems in CAT. Li et al. (2023) utilize the Deep Q-Network (DQN) to represent the action-value function $Q_{w}(s,q)$, representing the value of choosing question $q$ in state $s$, and $w$ denotes its parameter of the network’s fully connected layer. The most suitable question is selected according to the policy:
(9) $\pi^{*}(q|s)=\arg\max_{q\in\mathcal{Q}}{Q}_{w}(s,q).$
For modeling the complex interactions between examinees and questions during the practical test, Zhuang et al. (2022b) recently design a transformer-based Q-Network, which can identify and extract the perturbations in examinee performance (e.g., guess and slip factors). Stochastic Shortest Path Formulation. Furthermore, CAT can be defined as a Stochastic Shortest Path (SSP) problem (Bertsekas and Tsitsiklis, 1991), which is a special case of MDP. In an SSP, the objective is to find the shortest path (i.e., the minimum test step) from a given initial state $s_{0}$ to goal states. In CAT, the goal state typically represents the completion of the test or the attainment of a predetermined level of proficiency estimation precision. The approach by Gilavert and Freire (2022) uses Linear Programming to find the optimal testing policy $\pi^{*}$, treating the problem like a flow network where each state must have balanced inflow and outflow, except for the start and end points. It denotes variables $x_{s,a}$ as the expected accumulated occurrence frequency for every pair (state $s\in S$, action $q\in\mathcal{Q}$), and equalizes $in(s)$ and $out(s)$ flow model for every state $s$. The flow into a state $s$ is the sum of the expected frequencies of all actions in all other states $s^{\prime}$ that lead to $s$: ${in}(s)=\sum_{s^{\prime},q}x_{s^{\prime},q}P(s|s^{\prime},q)$. The flow out of a state $s$ is the sum of the expected frequencies of all actions in state $s$: ${out}(s)=\sum_{q}x_{s,q}$. The objective function is to maximize the total expected reward $r$, which is the sum of the expected frequencies times the immediate rewards $r(s,q)$ for all state-action pairs: $\min_{x_{s,q}}\sum_{s\in S,q\in\mathcal{Q}}{x_{s,q}r(s,q)}$. Thus the optimal policy $\pi^{*}$ can be obtained by:
(10) $\pi^{*}(q|s)=\frac{x_{s,q}}{\sum_{q^{\prime}\in\mathcal{Q}}x_{s,q^{\prime}}}.$
Partial-Observable MDP Formulation. Partial-Observable MDP (POMDP) is a generalization of MDP to handle situations where the agent does not fully observe the environment (Hoerger and Kurniawati, 2021). Previously, it assumes that the proficiency of an examinee can be perfectly inferred from their previous responses, then CAT can be modeled as an MDP, where the proficiency estimate can be the states. However, in a more general scenario, the proficiency cannot be perfectly inferred from their previous responses. Because there is always some uncertainty involved (Zhuang et al., 2022a; Drousiotis et al., 2021):
  1. (1)
    Perturbations in Responses: The assessment process is subject to various perturbations stemming from examinee behavior and the testing environment. For example, an examinee might guess an answer correctly or misunderstand a question they would typically answer correctly. This variability introduces uncertainty and error in estimation.
  2. (2)
    Limited Information: Particularly at the onset of an assessment, the test has only sampled a small portion of the examinee’s knowledge through limited questions. Consequently, the responses may not be entirely indicative of the examinee’s comprehensive proficiency.
  3. (3)
    Complexity of “proficiency”: Proficiency isn’t a simple attribute. It can encompass various aspects like understanding, analysis, and application of all the knowledge concepts. Responses to a constrained set of questions may fall short of capturing the full breadth of these dimensions, potentially leading to an incomplete portrayal of an examinee’s proficiency.
To this end, many works (Nurakhmetov, 2019; Gilavert and Freire, 2022; Chen et al., 2018) model CAT as a POMDP. Compared with MDP, the POMDP model has two additional elements. $O$: A set of observations; $Z$: Observation probabilities. $Z(o|s^{\prime},q)$ is the probability of making observation $o$ after selecting question $q$ and transitioning to state $s^{\prime}$. The states can still represent the proficiency of the examinees, but now these states are not fully observable. Instead, these works maintain a probability distribution $b(s)$, called the belief state, over the possible proficiencies of the examinee given his/her previous responses. The belief state is typically updated using Bayes’ rule. When the agent select action (question) $q$ in belief state $b$ and makes observation $o$, it updates its belief state to $b^{\prime}(s^{\prime})$: $b^{\prime}(s^{\prime})=\eta{Z(o|s^{\prime},q)\sum_{s\in S}P(s^{\prime}|s,q)b(s)}$. where $\eta$ is a normalizing constant. POMDPs can be solved by many algorithms, such as Grid-based algorithms (Hoerger and Kurniawati, 2021), Monte Carlo tree search (Schwartz et al., 2022). However, due to its partial observability of the environment, POMDPs are more challenging to solve than MDPs.
5.4. Meta Learning Algorithms
Another data-driven machine learning approach that can address this complex CAT problem is meta-learning (Finn et al., 2017): It involves training a model on various tasks to acquire cross-task knowledge or learn how to learn efficiently. Specifically, the base-learner is trained across a variety of related tasks, allowing it to gather cross-task insights and general knowledge about how to learn efficiently. Then, the meta-learner leverages this knowledge to swiftly adapt to new, unseen tasks (Huisman et al., 2021). In CAT, each examinee’s testing process can be seen as a task because it involves selecting appropriate test questions based on his/her own proficiency level. The selection algorithm can be regarded as a form of general knowledge because it represents the accumulated knowledge and experience gained from a diverse set of examinees (Figure 7(right)). This knowledge can include the best policy for question selection, information about the characteristics of different test questions, the examinee proficiency prior, etc. By learning from these diverse examinees (tasks) in the large-scale response dataset, it can acquire a good question selection that can adapt to individual examinees.
@@ 5.4. Meta Learning Algorithms
Figure 7. The overall Meta Learning framework of CAT, and this figure is adapted from (Ghosh and Lan, 2021). The objective is to optimize the selection algorithm $\pi$ by exploring the large-scale examinee response data.
@@
Bi-Level Optimization. One particular class of meta-learning methods that has shown promise is bi-level optimization, which involves optimizing both the global model’s parameters and the task-specific parameters simultaneously. This approach allows the model to learn how to adapt its learning process to different tasks, resulting in improved performance and generalization. Ghosh and Lan (2021) proposes a bi-level optimization framework for CAT (BOBCAT) to directly learn the data-driven selection algorithm $\pi$. Specifically: let $N$ denote the number of examinees in the response dataset for training $\pi$. The responses of each examinee $i$ are randomly divided into a support set $\mathcal{D}_{s}^{i}$ and a query set $\mathcal{D}_{u}^{i}$, where $\pi$ sequentially select a total of $t$ questions $\{q_{1},...,q_{t}\}$ from $\mathcal{D}_{s}^{i}$, observe their responses, and predict their response on the held-out query set $\mathcal{D}_{u}^{i}$. The global knowledge (i.e., selection algorithm $\pi$ and global parameters $\gamma$) is redefined as the objective of bi-level optimization:
(11) $\min_{\pi,\gamma}\frac{1}{N}\sum_{i=1}^{N}{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{l(y,f(q,\hat{\theta}_{i}))}},$
(12) $\mathrm{s.t.}\;\;\hat{\theta}_{i}=\mathop{\arg\min}_{\theta_{i}}{\sum_{(q,y)\in\mathcal{D}_{s}^{i}}{l\left(y,f\left(q,\theta_{i}\right)\right)}},$
(13) $\mathrm{where}\quad q_{t+1}\sim\pi\left(q|q_{1},y_{i(1)},...,q_{t},y_{i(t)}\right)\in\mathcal{D}_{s}^{i}.$
@@
Figure 7(left) shows the overall meta learning framework. In the inner-level (Eq.(12)), the question in the support set $\mathcal{D}_{s}^{i}$ for examinee $i$ is sequentially selected by $\pi$, according to his/her previous responses; then binary cross-entropy loss $l(\cdot)$ on $\mathcal{D}_{s}^{i}$ is minimized for estimating the proficiency $\hat{\theta}_{i}$ for the outer-level. In the outer-level (Eq.(11)), the loss of the estimate $\hat{\theta}_{i}$ on the query set $\mathcal{D}_{u}^{i}$ is minimized to learn the selection algorithm $\pi$ and the global parameters $\gamma$. The global parameters $\gamma$ refer to the question characteristics and the examinee ability prior. The algorithm $\pi$ is also model-agnostic. It could be adapted to the given CDM ($f$) automatically by optimizing this problem for efficient selection. Once the question selection algorithm is trained, its parameters do not update during the CAT process and adaptively selects the next item based on previous response behaviors. Based on the above bi-level definition, there have been increasing efforts to improve upon it. Ma et al. (2023) propose a flexible optimization framework Decoupled Learning CAT (DL-CAT). The original BOBCAT obtains the parameters of two modules (i.e., examinee proficiency estimation and question selection algorithm) through coupled inner and outer optimizations, i.e., the result of the outer optimization model is used to measure the quality of the inner. DL-CAT devises a ground-truth construction strategy, and a pairwise loss function, allowing these two models to be trained independently; Feng et al. (2023) introduces a constrained version of BOBCAT to address the question exposure and test overlap issues. Yu et al. (2023) recently introduce the collaborative information of examinees in optimizing this bi-level problem, achieving fast convergence of proficiency estimation. Meta-Learning vs Reinforcement Learning. In CAT, meta-learning can be seen as a higher-level learning process that learns how to adapt a general strategy for question selection to specific examinees based on their responses. Actually, this meta-learning CAT can be reframed as a RL problem. Zhuang et al. (2022b) propose NCAT which transforms the meta-learning problem in CAT into a RL problem. Because the test may stop at any step according to different stopping rules, NCAT simplifies the original objective (Eq(11)) in meta-learning and sums all the test steps to minimize the loss:
@@
(14) $\min_{\pi}\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{l(y,f(q,\hat{\theta}_{i}^{t}))}}$$\triangleq$ $\mathop{\max}\limits_{\pi}\mathbb{E}_{i\sim\pi}\left[\sum_{t=1}^{T}{-{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{l(y,f(q,\hat{\theta}_{i}^{t}))}}}\right]$
$=$ $\max_{\pi}\mathbb{E}_{i\sim\pi}\left[\sum_{t=1}^{T}{-L({\hat{\theta}_{i}^{t}};\mathcal{D}_{u}^{i})}\right],$
@@
where $\hat{\theta}_{i}^{t}=\mathop{\arg\min}_{\theta_{i}}{\sum_{(q,y)\in\mathcal{D}_{s}^{i(t)}}{l\left(y,f\left(q,\theta_{i}\right)\right)}}$ and $\mathcal{D}_{s}^{i(t)}=\{q_{1},y_{i(1)},...,q_{t},y_{i(t)}\}$. Thus, the bi-level optimization is transformed into maximizing the expected cumulative reward (i.e., $-L({\hat{\theta}_{i}^{t}};\mathcal{D}_{u}^{i})$) in RL settings, where the reward is the negative loss of the estimated proficiency of examinee $i$ on the query set at step $t$. Other RL elements (e.g., state) in NCAT are the same in Section 5.3. Recently, Wang et al. (2023a) further design a Multi-Objective RL framework of CAT. It contains three objectives in the reward function, which can improve the prediction accuracy, increase the concept diversity and reduce the question exposure respectively.
Discussion 4.
The aforementioned data-driven machine learning approaches, i.e., Reinforcement Learning and Meta Learning, are capable of directly learning from data to optimize question selection algorithms (policy). By training a policy $\pi$ on extensive response datasets and selecting $T$ questions, they strive towards achieving the ultimate objective of CAT: $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|$. They can analyze learning behaviors, performance responses, and preferences to reveal underlying patterns and correlations. However, the training overhead for these methods is substantial, and there is also the potential risk of introducing biases present in the dataset.
5.5. Subset Selection Algorithms

### 09 Mar 2026 v3

@@ 5.1 Statistical Algorithms
Figure 3: The Active Learning Framework and the relationship/correspondence between each component of Active Learning and those of CAT.
5.2 Active Learning Algorithms
To design selection algorithms that are effective across different measurement models i.e., model-agnostic, researchers explore a general machine learning technique for data selection: Active Learning [85]. Active Learning is to actively choose some valuable data, thus can train better models with less data. This technique has improved data efficiency in numerous learning tasks [72].
As shown in Figure 3, active learning operates in cycles: where a selection algorithm iteratively chooses unlabeled samples based on the model’s current performance and queries a human annotator for labels. This process augments limited labeled data to improve model performance. The core challenge lies in designing effective sample selection algorithms, typically based on two criteria: informativeness, selecting samples that reduce model uncertainty [179], and representativeness, selecting samples that reflect the overall data distribution [56, 93], or a combination of both [169]. Active learning shares a similar structure with CAT. Here, the measurement model plays the role of the learning model, the question selection corresponds to sample selection, and examinee responses serve as annotations. The goal is to estimate proficiency using as few questions as possible. This model-agnostic perspective avoids reliance on specific measurement model assumptions. Bi et al. [16] propose MAAT, a model-agnostic adaptive testing framework that evaluates the change (analogous to a gradient) in proficiency estimates after each response:
$j_{t+1}=\arg\max_{q_{j}\in\mathcal{Q}}\mathbb{E}_{y_{j}}\left\|\nabla_{\theta}L(\mathcal{D}_{1:t}\cup\{(q_{j},y_{j})\}|\theta)\right\|.$ (9)
Since the true responses to candidate questions are not available during selections, it computes the expected gradient norm with respect to the response label $y$ for each candidate question. This expectation quantifies the potential impact of each question on the proficiency estimation. The intuition behind this framework is that it prefers questions that are likely to most influence the proficiency estimation (i.e., have the greatest impact on its parameters). Notably, this approach places no restriction on the specific type of measurement model, as long as it supports gradient-based optimization.
Discussion: With the rapid development of intelligent testing platforms (ranging from human’s online testing systems to AI model’s evaluation leaderboards), large-scale examinee response data has been accumulated. However, such data cannot be effectively leveraged by the above rules-based approaches (i.e., statistical algorithms and Active Learning algorithms) [57, 97]. In contrast, recent data-driven approaches based on Reinforcement Learning (Section 5.3) and Meta-Learning (Section 5.4) have gained increasing attention. They automatically learn/optimize effective selection algorithms from large-scale response data without relying on manually defined heuristics or rules, and have demonstrated superior performance.
5.3 Reinforcement Learning Algorithms
@@
Reinforcement learning (RL), a subfield of machine learning, is a powerful approach that enables an agent to learn how to make optimal decisions automatically [148]. It has been successfully applied in various domains, including robotics, autonomous vehicles, education, and healthcare [6, 83]. In RL, an agent interacts with an environment and receives feedback in the form of rewards or penalties based on its actions. The goal is to learn a policy $\pi$, which can maximize the long-term cumulative reward. As shown in Figure 4, the policy can be learned by exploring the environment and learning from its consequences of actions. In essence, researchers in CAT utilize RL methodologies to address a question: Can the selection algorithm (policy) be automatically learned and optimized from data or examinee interactions, thus circumventing the necessity for expert intervention?
Markov Decision Process Formulation. The interaction between agent and environment can be viewed as a Markov Decision Process (MDP) [50]. Specifically, at each step, the agent observes current environment’s state ($s$), and interacts with the environment by selecting its actions ($a$). Simultaneously, the agent receives a reward ($r$) from these interactions, influencing or changing the current state of the environment. The objective is to select a best sequence of actions, resulting in the highest cumulative reward ($\sum_{t}r_{t}$). Therefore, most RL problems are formally described as estimating the optimality of the agent’s behavior in a given state (value-based methods [148]) or the optimality of the action policy itself (policy-based methods [2]) or the hybrid approaches [116]. The overall RL framework for CAT is illustrated in Figure 4. We formulate the CAT problem as an MDP, where the key RL components in the testing system are defined below:
@@
  • State: A state $s_{t}\in\mathcal{S}$ represents the current condition or situation at each test step $t$. It captures relevant information about examinee and the CAT system. Generally, the state includes the examinee’s previous response sequence (or a latent vector to represent the current proficiency estimate [95]) and the candidate questions in the question bank [189, 121]: $s_{t}=(\{q_{1},y_{1},...,q_{t},y_{t}\},\mathcal{Q})$ [^1] This aligns with the labeled data (answered questions), and the unlabeled data (question bank) in active learning (Section 5.2).
@@
  • Reward: A reward $r$ is a scalar feedback that the CAT receives after selecting a question for the examinee. To achieve CAT’s goal in Definition 1, the reward function can be defined as the accuracy of proficiency estimation at each step [^2] As the true value is often unobtainable, it is commonly derived through simulation experiments.[108, 121, 168], i.e., $\|\hat{\theta}^{t}-\theta_{0}\|$, or the performance prediction loss of $\hat{\theta}^{t}$ on the held-out response data $\mathcal{D}$ [189, 144], i.e., $L(\mathcal{D}|\hat{\theta}^{t})$. This reward signal is pivotal in guiding the policy $\pi$ to select the best-fitting question that can reduce the estimation error.
@@
Figure 4: The overall Reinforcement Learning framework of CAT. The objective is to optimize the selection algorithm $\pi$ (i.e., policy) by exploring the large-scale examinee response data (i.e., environment).
Recently, with the advancements in deep learning, an increasing number of studies are leveraging Deep Reinforcement Learning to tackle the MDP problems in CAT. Li et al. [96] utilize the Deep Q-Network to represent the action-value function $Q_{w}(s,q)$, representing the value of choosing question $q$ in state $s$, and $w$ denotes its parameter of the network’s fully connected layer. The most suitable question is selected according to the policy:
$\pi^{*}(q|s)=\arg\max_{q\in\mathcal{Q}}{Q}_{w}(s,q).$ (10)
To further capture the complex interactions between examinees and questions in practical testing scenarios, a Transformer-based Q-Network named NCAT [189] has been proposed. NCAT incorporates multiple functional modules, including a Double-Channel Performance Learning module that independently captures diverse aspects of examinee performance, and a Contradiction Learning module that identifies and extracts inconsistencies in examinee behavior, such as guessing and slipping.
Stochastic Shortest Path Formulation. Furthermore, CAT can be defined as a Stochastic Shortest Path (SSP) problem [14], which is a special case of MDP. In an SSP, the objective is to find the shortest path (i.e., the minimum test step) from a given initial state $s_{0}$ to goal states. In CAT, the goal state typically represents the completion of the test or the attainment of a predetermined level of proficiency estimation precision. Gilavert et al. [61] use Linear Programming to find the optimal testing policy $\pi^{*}$, treating CAT like a flow network where each state must have balanced inflow and outflow (except for the start and end points). It denotes variables $x_{s,a}$ as the expected accumulated occurrence frequency for every pair (state $s\in\mathcal{S}$, qu $q\in\mathcal{Q}$), and equalizes $in(s)$ and $out(s)$ flow model for every state $s$. The flow into a state $s$ is the sum of the expected frequencies of all actions in all other states $s^{\prime}$ that lead to $s$: ${in}(s)=\sum_{s^{\prime},q}x_{s^{\prime},q}P(s|s^{\prime},q)$. The flow out of a state $s$ is the sum of the expected frequencies of all actions in state $s$: ${out}(s)=\sum_{q}x_{s,q}$. The objective function is to maximize the total expected reward $r$, which is the sum of the expected frequencies times the immediate rewards $r(s,q)$ for all state-action pairs: $\min_{x_{s,q}}\sum_{s\in\mathcal{S},q\in\mathcal{Q}}{x_{s,q}r(s,q)}$. Thus the optimal question selection policy $\pi^{*}$ can be obtained by:
$\pi^{*}(q|s)=\frac{x_{s,q}}{\sum_{q^{\prime}\in\mathcal{Q}}x_{s,q^{\prime}}}.$ (11)
Partial-Observable MDP Formulation. Partial-Observable MDP (POMDP) extends the standard MDP framework to settings where the environment is only partially observable [46]. Traditional CAT models often assume that an examinee’s proficiency can be fully inferred from previous responses, allowing it to be treated as an MDP with the proficiency estimate as the state. However, in practice, proficiency cannot be perfectly inferred due to some inherent uncertainty [188, 47].
To this end, many works [121, 61, 29] model CAT as a POMDP. Compared with MDP, the POMDP model has two additional elements. $O$: A set of observations; $Z$: Observation probabilities. $Z(o|s^{\prime},q)$ is the probability of making observation $o$ after selecting question $q$ and transitioning to state $s^{\prime}$. While the underlying state (proficiency) remains, it is not fully observable. Instead, these methods maintains a belief state $b(s)$, a probability distribution over possible proficiencies, which is updated via Bayes’ rule: When the agent select action (question) $q$ in belief state $b$ and makes observation $o$, it updates its belief state to $b^{\prime}(s^{\prime})$: $b^{\prime}(s^{\prime})=\eta{Z(o|s^{\prime},q)\sum_{s\in\mathcal{S}}P(s^{\prime}|s,q)b(s)}$. where $\eta$ is a normalizing constant. POMDPs can be solved by many algorithms, such as Grid-based algorithms [69], Monte Carlo tree search [140]. However, due to its partial observability of the environment, POMDPs are more challenging to solve than MDPs.
5.4 Meta Learning Algorithms
Another data-driven machine learning approach that can address this complex CAT problem is meta-learning [53]: It involves training a model on various tasks to acquire cross-task knowledge or learn how to learn efficiently. Specifically, the base-learner is trained across a variety of related tasks, allowing it to gather cross-task insights and general knowledge about how to learn efficiently. Then, the meta-learner leverages this knowledge to swiftly adapt to new, unseen tasks [177]. In CAT, each examinee’s testing process can be seen as a task because it involves selecting appropriate test questions based on the proficiency level. The selection algorithm can be regarded as a form of general knowledge because it represents the accumulated knowledge and experience gained from a diverse set of examinees (Figure 5). This knowledge can include the best policy for question selection, information about the characteristics of different test questions, the examinee proficiency prior, etc. By learning from these diverse examinees (tasks) in the large-scale response dataset, it can acquire a good question selection that can adapt to individual examinees.
Figure 5: The overall Meta Learning framework of CAT, and this figure is adapted from [57]. The objective is to optimize the selection algorithm $\pi$ by exploring the large-scale examinee response data.
@@ 5.4 Meta Learning Algorithms
Bi-Level Optimization. Bi-Level optimization is a classical meta-learning approach commonly applied in CAT. It decomposes the learning process into two nested levels: an inner level that adapts to individual examinees and an outer level that learn general knowledge. Ghosh et al. [57] propose a bi-level optimization framework for CAT (BOBCAT) to directly learn the data-driven selection algorithm $\pi$. Specifically: let $N$ denote the number of examinees in the response dataset for training $\pi$. The responses of each examinee $i$ are randomly divided into a support set $\mathcal{D}_{s}^{i}$ and a query set $\mathcal{D}_{u}^{i}$, where $\pi$ sequentially select a total of $t$ questions $\{q_{1},...,q_{t}\}$ from $\mathcal{D}_{s}^{i}$, observe their responses, and predict their response on the held-out query set $\mathcal{D}_{u}^{i}$. The global knowledge (i.e., selection algorithm $\pi$ and global parameters $\gamma$) is redefined as the objective of bi-level optimization:
$\min_{\pi,\gamma}\frac{1}{N}\sum_{i=1}^{N}{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{\ell(y,f(q,\hat{\theta}_{i}))}},$ (12)
$\mathrm{s.t.}\;\;\hat{\theta}_{i}=\mathop{\arg\min}_{\theta_{i}}{\sum_{(q,y)\in\mathcal{D}_{s}^{i}}{\ell\left(y,f\left(q,\theta_{i}\right)\right)}},$ (13)
$\mathrm{where}\quad q_{t+1}\sim\pi\left(q|q_{1},y_{i(1)},...,q_{t},y_{i(t)}\right)\in\mathcal{D}_{s}^{i}.$ (14)
Figure 5 shows the overall meta learning framework. In the inner-level (Eq.(13)), the question in the support set $\mathcal{D}_{s}^{i}$ for examinee $i$ is sequentially selected by $\pi$, according to the previous responses; then binary cross-entropy loss $\ell(\cdot)$ on $\mathcal{D}_{s}^{i}$ is minimized for estimating the proficiency $\hat{\theta}_{i}$ for the outer-level. In the outer-level (Eq.(12)), the loss of the estimate $\hat{\theta}_{i}$ on the query set $\mathcal{D}_{u}^{i}$ is minimized to learn the selection algorithm $\pi$ and the global parameters $\gamma$ (e.g., question characteristics). The algorithm $\pi$ is also model-agnostic. It could be adapted to the given measurement model ($f$) automatically by optimizing this problem for efficient selection.
Through large-scale sampling and training, this framework learns to estimate and quantify the value of each question for different examinees and under varying contexts. Even for questions whose IDs do not appear in the training set, their value can be inferred from their characteristics via $\gamma$. Once the question selection algorithm is trained, its parameters do not update during the CAT process and adaptively select the next question based on previous response behaviors.
@@
Based on BOBCAT, there have been increasing efforts to improve upon it. Ma et al. [108] propose a flexible optimization framework Decoupled Learning CAT (DL-CAT). The original BOBCAT obtains the parameters of two modules (i.e., examinee proficiency estimation and question selection algorithm) through coupled inner and outer optimizations, i.e., the result of the outer optimization model is used to measure the quality of the inner. DL-CAT devises a ground-truth construction strategy, and a pairwise loss function, allowing these two models to be trained independently; Feng et al. [52] introduces a constrained version of BOBCAT to address the question exposure and test overlap issues. Yu et al. [180] recently introduce the collaborative information of examinees in optimizing this bi-level problem, achieving fast convergence of proficiency estimation.
Meta-Learning vs Reinforcement Learning. In CAT, meta-learning methods can be seen as a higher-level learning process that learns how to adapt a general strategy for question selection to specific examinees based on their responses. Actually, it can be reframed as an RL problem. Zhuang et al. [189] propose NCAT to transform the meta-learning problem in CAT into an RL problem. Because the test may stop at any step according to different stopping rules, NCAT simplifies the original objective (Eq(12)) and sums all the steps to minimize the loss:
$\min_{\pi}\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{\ell(y,f(q,\hat{\theta}_{i}^{t}))}}$ (15)
@@
$\triangleq$ $\mathop{\max}\limits_{\pi}\mathbb{E}_{i\sim\pi}\left[\sum_{t=1}^{T}{-{\sum_{(q,y)\in\mathcal{D}_{u}^{i}}{\ell(y,f(q,\hat{\theta}_{i}^{t}))}}}\right]$
$=$ $\max_{\pi}\mathbb{E}_{i\sim\pi}\left[\sum_{t=1}^{T}{-L(\mathcal{D}_{u}^{i}|{\hat{\theta}_{i}^{t}})}\right],$
@@
where $\hat{\theta}_{i}^{t}=\mathop{\arg\min}_{\theta_{i}}{\sum_{(q,y)\in\mathcal{D}_{s}^{i(t)}}{\ell\left(y,f\left(q,\theta_{i}\right)\right)}}$ and $\mathcal{D}_{s}^{i(t)}=\{q_{1},y_{i(1)},...,q_{t},y_{i(t)}\}$. Thus, the bi-level optimization is transformed into maximizing the expected cumulative reward (i.e., $L(\mathcal{D}_{u}^{i}|{\hat{\theta}_{i}^{t}})$) in RL settings, where the reward is the negative loss of the estimated proficiency of examinee $i$ on the query set at step $t$. Recently, GMOCAT [168] has been proposed as a Multi-Objective RL framework. GMOCAT uses Graph Neural Networks to capture the complex relationships between questions and skills. It adopts an Actor-Critic architecture and incorporates three objectives into the reward function: (1) improving prediction accuracy, (2) enhancing concept (skill) diversity, and (3) reducing question exposure.
Discussion: The aforementioned data-driven machine learning approaches, i.e., Reinforcement Learning and Meta Learning, are capable of uncovering latent patterns and correlations from data, and directly optimizing question selection policies. By fitting to large-scale data, they can approximate the ultimate goal of CAT. However, potential issues such as data bias, model overfitting, and high training overhead should not be overlooked.
Figure 6: Illustration of the subset optimization problem, adapted from [191]: Selecting subset $S$ to cover the bank $Q$. Rectangles represent different questions, with $w(i,j)$ measuring the similarity of question pair.
5.5 Subset Selection Algorithms